OpenStudy (anonymous):

To get $3000 after 2.75 years at 5% compounded daily. the present value of $3000 is $ ______

jimthompson5910 (jim_thompson5910):

You use the formula \[\Large FV = PV\left(1+\frac{r}{n}\right)^{nt}\]

jimthompson5910 (jim_thompson5910):

PV = present value FV = future value r = interest rate as a decimal n = compounding frequency t = time in years

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Here, just simplify this:\[3000=PV(1+\frac{ .05 }{ 365 })^{2.75\times365}\]